OpenStudy (anonymous):

Please help, exam in 30 min, need to know how to do this question. (Charles Law) The temp in my fridge is about 4Celsius, if i place a balloon in my fridge that initially has a temp of 22Celsius and a volume of 0.5 L, what will be the volume of the balloon whenn it is fully cooled by fridge???? Please, if someone could show me how to do this for my exam, it would be extremelly appreciated!!!

OpenStudy (rogue):

Do you want me to tell you how to do it or just post the work & solution?

OpenStudy (anonymous):

both i suppose, but really just need to find how to work out the equation to reach the answer

OpenStudy (anonymous):

:)

OpenStudy (rogue):

Step 1: Convert your temperatures from Celsius to Kelvin. C + 273 = K 4 C + 273 = 277 k 22 C + 273 = 295 k Step 2: Use Charles's Law to find your new volume (V2). Your initial conditions are V1 and T1, which is 0.5 L and 295 k. Your final conditions are V2, which you are going to find out, and T2, which is 277 k. \[Charles's .Law : \frac {V_1}{V_2} = \frac {T_1}{T_2}\]\[\frac {0.5 L}{V_2} = \frac {295 k}{277 k}\]\[V_2 = \frac {277 k * 0.5 L}{295 k} \approx 0.47 L\]

OpenStudy (anonymous):

that is awesome, i totally understand it now. time for my exam. O.O thanks a bunch very awesome. peace.

OpenStudy (rogue):

Good luck! :)